OpenStudy (anonymous):

Calculate the density of an unknown material that has a mass of 352.49 g and a volume of 148.5 cm3.

OpenStudy (anonymous):

do baby lizards bite?

OpenStudy (asib1214):

To fidn the density, remember the formula D= m/v 148.5cm^3 =5.29L D = 352.49g/5.29l = 66.63

OpenStudy (anonymous):

Okay. Thanks c: